Ali, who has a mass of 43.0 kg, slides onto a handrail in his grandfather's house. If the handrail has an angle of 35.0 to the level of the horizontal, what is the vertical force between Ali and the handrail?


1
Expert's answer
2021-02-15T00:54:07-0500
"F_N=mgcos\theta,""F_N=43\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot cos35^{\circ}=345\ N."